# Accurate Screening of Functional Materials with Machine-Learning Potential and Transfer-Learned Regressions: Heusler Alloy Benchmark

### v1

@@ 4.3 Computational methods
The ML-HTP candidate list and DFT validation results are included in the Supplementary Materials. The complete set of all screened compounds, along with ML-predicted properties, will be made available through the HeuslerDB database at https://www.nims.go.jp/group/spintheory/ .
@@
The underlying code for this study is not publicly available but may be made available to qualified researchers on reasonable request from the corresponding author.
@@ Acknowledgments
This study was supported by MEXT Program: Data Creation and Utilization-Type Material Research and Development Project (Digital Transformation Initiative Center for Magnetic Materials) Grant Number JPMXP1122715503 and as “Program for Promoting Researches on the Supercomputer Fugaku” (Data-Driven Research Methods Development and Materials Innovation Led by Computational Materials Science, JPMXP1020230327). This study used computational resources of supercomputer Fugaku provided by the RIKEN Center for Computational Science (Project ID: hp240223), the computer resources provided by ISSP, U-Tokyo under the program of SCCMS, and the computer resources at NIMS Numerical Materials Simulator.

### v2

@@ 4.3 Computational cost
The ML-HTP candidate list and DFT validation results are included in the Supplementary Information as Tables S3 and S4. The complete set of all screened compounds, along with ML-predicted properties, will be made available through the HeuslerDB database at https://www.nims.go.jp/group/spintheory/.
@@
The developed packages MLIP-HOT and MLIP-FTL will be made available through the Spin Theory Group GitHub repository at https://github.com/nims-spin-theory and our group website at https://www.nims.go.jp/group/spintheory/.
Funding
This study was supported by MEXT Program: Data Creation and Utilization-Type Material Research and Development Project (Digital Transformation Initiative Center for Magnetic Materials) Grant Number JPMXP1122715503 and as “Program for Promoting Researches on the Supercomputer Fugaku” (Data-Driven Research Methods Development and Materials Innovation Led by Computational Materials Science, JPMXP1020230327).
@@ Acknowledgments
This study used computational resources of supercomputer Fugaku provided by the RIKEN Center for Computational Science (Project ID: hp240223), the computer resources provided by ISSP, U-Tokyo under the program of SCCMS, and the computer resources at NIMS Numerical Materials Simulator.